# Role-Selection Game in Block Production under Proposer-Builder Separation

### v1

@@ II-B Modeling of Co-evolution
In our modeling, we employ genetic algorithms to capture the co-evolution process of agents. Each agent begins with a randomly generated population of 20 strategies. The block building strategy comprises a 5-digit binary sequence referred to as the ”chromosome,” indicating the relevant rebate ratio $\alpha_{j}$. As shown in Fig. 2, for bundle sharing, $S_{i,r,k}$ signifies the $k$th strategy of searcher $i$ in the $r$th generation. Each 5-digit segment within a chromosome corresponds to a parameter value that regulates the bidding behaviors of bundle sharing strategy. The bid ratio is computed using a modified sigmoid function
@@
where $\gamma_{i,1}\in[1,5]$ measures the sensitivity of searchers’ bidding inclinations towards builders with elevated rebate ratio, whereas $\gamma_{i,2}\in[0,4]$ determines the overall scale of the bid ratio. The adoption of the modified sigmoid function ensures that the bid ratio $\beta_{ij}$ remains within the interval $[0,1]$ and is monotonically increasing in relation to $\alpha_{j}$. The largest decimal number that a 5-digit binary number can represent is $31$. Taking $S_{i,r,1}$ ”0010101001” as an example, the first 5-digit binary segment ”00101” can be translated to the decimal number number $51\cdot 2^{2}+1\cdot 2^{0}=$ 5. Thus, the parameter $\gamma_{i,1}$ can be derived from ”00101” as $1+(5/31)\cdot(5-1)=21/31$. Similarly, the second 5-digit binary ”01001” corresponds to the parameter $\gamma_{i,2}=0+(9/31)\cdot(4-0)=36/31$ suggesting that $S_{i,r,1}$ ”0010101001” corresponds to the parameter pair $(\gamma_{i,1},\gamma_{i,2})=(21/31,36/31)$. The rebate ratio $\alpha_{j}$ for block building is directly obtained through a mapping that converts a 5-digit binary number into a decimal value within the interval $[0,1]$.
Figure 2: Searcher $i$’s population of strategies in the $r$th generation.
@@
Step 1: A private value is randomly generated for each agent, representing an opportunity available to them, alongside the random creation of an interaction matrix. In our subsequent experiments, the private values adhere to an exponential distribution with a default parameter of 10 (mean value of 0.1) [^2] Given a fixed mean (or rate parameter), the exponential distribution is the exclusive distribution that fulfills the maximum entropy requirement, thereby introducing the fewest extra assumptions.. For simplicity, the interaction matrix is constructed according to a two-point distribution $\mathbb{P}(\varphi_{ij}=\varphi_{ji}=-1)=p_{C}$ and $\mathbb{P}(\varphi_{ij}=\varphi_{ji}=0)=1-p_{C}$. This implies that for any two agents, they completely conflict with a probability of $p_{C}$, or they operate independently with a probability of $1-p_{C}$  [^3] As pointed in [17], order flow providers possessing complementary flows are incentivized to integrate in order to collaboratively capture greater value, which consequently leads to a form of interaction between bundles that is predominantly characterized by conflict..
@@
$\mathbb{P}(S_{i,r,k}~{}\text{is selected})=\frac{e^{\frac{f(S_{i,r,l})}{T}}}{\sum_{k}e^{\frac{f(S_{i,r,k})}{T}}},$
where $T$ controls the trade-off between exploration and exploitation. Through experimentation, we found that $T=2$ is an appropriate value. At the start of the simulation, agents were randomly assigned a population of 20 strategies. Agents will transform the selected strategy (”chromosome”) into executable actions (bid ratios or rebate ratios). For builders (block building), they will determine their rebate ratio $\alpha_{j}$, while for searchers (bundle sharing), they will establish their bid ratio $\bm{\beta}_{i}$.
Step 3: Based on the individual strategies of the agents, the simulation of the block building process unfolds: searchers send bundles to builders, who utilize the bundles received from searchers along with their own bundles to construct the valuable blocks. We employed a greedy algorithm (Algorithm 1) to simulate the merging of bundles into a block  [^4] This study focuses on conditions of sufficient block space, wherein the total quantity of bundles does not exceed the capacity of the block. Nevertheless, our findings can be readily adjusted to address cases with capacity limitations on blocks, where a block can accommodate a maximum of $L$ bundles.. Builders participate in the block building auction to calculate the final rewards received by both builders and searchers from the winning block. Update the strategy fitness based on the strategies selected by the agents in this round.
@@
Step 4: Agents optimize their strategy repository through genetic algorithms with a trigger probability of $0.01$. Initially, a subset of strategies is removed from the repository based on a specified elimination ratio (configured as $0.5$ in the simulation). Grounded in the ”chromosomes,” genetic algorithms integrate selection, crossover, and mutation processes until the strategy repository is replenished back to its upper limit of $20$, as detailed below.
@@
  • Crossover: Two parent strategies are combined to create two offspring strategies by exchanging segments of the ”chromosome” representations of the parent strategies. The fitness of offspring strategies is calculated as the mean of the fitness values of their parent strategies.
  • Mutation: For every bit in the binary chromosomes of each offspring produced by crossover, a flip ($0\rightarrow 1$ and $1\rightarrow 0$) is performed with a specified mutation rate of $0.01$.
III Results
In this section, we present the results of agent-based simulations. First, we examine the outcomes of the co-evolution of agents based on genetic algorithms. Next, we analyze the impact of bundle interactions on system evolution and the respective payoffs. Finally, we evaluate and predict the role selection of participants in the block building market using empirical game-theoretic analysis.
III-A Co-evolution of Agents’ Strategies
Consider a model comprising 10 builders and 10 searchers, where the probability of conflict $p_{C}=0.8$. We run the model for 10,000 iterations, which yields 10,000 blocks; each iteration is referred to as a generation.
First, we examine the convergence of agents’ strategies using genetic algorithms. For a single builder, within a single generation, there are 20 strategies maintained in the strategy repository, where each strategy of rebate ratio, encoded as 5-bit binary codes, can be transformed into decimal values (integers ranging from 0 to 31). We compute the coefficient of variation (CoV) of these decimal values, defined as the ratio of the standard deviation to the mean value. Therefore, the convergence of builders’ rebate ratios can be indicated by a reduction in the CoV. Likewise, the convergence of searchers’ behavioral parameters $\gamma_{1}$ and $\gamma_{2}$ can be represented by the decline of CoV as well. As agents employing the same meta-strategy are homogeneous, we calculate the average CoV for the rebate ratios $\alpha$ of all builders and the average CoV for all searchers corresponding to parameters $\gamma_{1}$ and $\gamma_{2}$ in each generation. This allows us to derive the evolution of CoV as illustrated in Fig. 3. From Fig. 3, we can observe that CoV of $\alpha$, $\gamma_{1}$ and $\gamma_{2}$ all exhibit a declining trend, decreasing from an initial value of 0.6 to values of 0.08, 0.28, and 0.25, respectively, by the 10000th generation. The convergence process of the CoV of $\gamma_{1}$ is notably more stable, whereas the decline of the CoV of $\alpha$ and CoV of $\gamma_{2}$ are characterized by significant periodic fluctuations. This suggests that the sensitivity of searchers’ bidding inclinations concerning builders with higher rebate ratios is likely to exhibit stability. In contrast, searchers experience a cycle of divergence and convergence while optimizing the overall scale of the bid ratio, whereas builders possess a more dispersed distribution of rebate ratio strategies.
(a) Decline of CoV
@@ III-A Co-evolution of Agents’ Strategies
To examine the evolutionary process of agents’ behaviors, we calculated the cumulative fitness of all strategies within the agents’ strategy repository for each generation. Figs. 3, 3 and 3 depict the evolution of the cumulative fitness of $\gamma_{1}$, $\gamma_{2}$ and $\alpha$, respectively. The fitness reflects the performance of the strategy and the agents’ propensity to utilize it. On one hand, we observed a general differentiation in strategy fitness, with a small subset of strategies exhibiting high fitness, represented by the lighter shades in the figure, clustering around certain values, while the fitness of the remaining strategies approaches zero, consistent with the overall decline in CoV. Simultaneously, high-fitness strategies are not completely concentrated around a specific value, indicating a certain degree of diversity in the agents’ strategies. Under the current settings, we found that, overall, searchers tend to engage in differentiated bidding, specifically setting higher bid ratios towards builders that announce higher rebate ratios to share more value, as demonstrated in Fig. 3. Meanwhile, from Fig. 3, searchers generally tend to opt for smaller values of $\gamma_{2}$ to enhance their overall bid ratios, revealing the competitive landscape among searchers influenced by $p_{C}=0.8$. Regarding the rebate ratios of builders, Fig. 3 illustrates that several exceedingly high rebate ratios correspond to high fitness, indicating intense competition among builders and relatively low returns for them; we will analyze agents’ profitability in the subsequent subsections.
@@
To provide a clearer illustration of the overall evolution of rebate ratios and bid ratios among the two types of agents within the system, we calculated the average values of bid ratios (for searchers) and rebate ratios (for builders) based on the actual actions taken in each generation, as depicted in Figs. 3 and 3, respectively. The calculation of the moving average reveals a sustained upward trend in the average bid ratio, as shown in Fig. 3, alongside the convergence of the average rebate ratio toward approximately 0.7, as indicated in Fig. 3.
Overall, through the examination of the evolutionary processes of agents, we find that our agent-based simulation, founded on genetic algorithms, has been successful. Within the PBS interactive environment we constructed, agents have autonomously learned strategies to adapt to their surroundings based on reinforcement learning principles, which involves a co-evolutionary process among agents. Ultimately, the strategies of the agents converge while retaining a certain degree of diversity.
III-B Impact of Bundle Interactions
@@
In this subsection, we focus on the impact of bundle interactions on the evolutionary outcomes of the system. During the simulation process, we simplify the interactions among bundles to include conflicts that arise between any two bundles with a specified probability. By performing 10 repetitions for each designated conflict probability within a system comprising 5 builders and 5 searchers, we statistically analyze the average bid ratio of searchers, the average rebate ratio of builders, and the profits of searchers and builders’ proposers after the system reaches a state of stability, yielding the results depicted in Figure 4. According to Fig. 4, it is evident that the average bid ratio and average reward results of searchers exhibit less volatility compared to those of builders. For searchers, an increase in the probability of conflicts significantly intensifies competition among them, elevating the overall level of their bid ratios while concurrently reducing their average returns, as depicted in Figs. 4 and 4.
The impact of the probability of conflicts on builders is relatively minor; however, it exhibits a more complex non-monotonic pattern. When the probability of conflicts is low, an increase in this probability encourages builders to raise their rebate ratios, a result of co-evolution: searchers correspondingly increase their bid ratios, thereby providing more value bundles to builders. In turn, builders enhance their rebate ratios to capitalize on the differentiated order flow from searchers. However, once the probability of conflicts surpasses a certain threshold (0.3 in Fig. 4), more frequent bundle conflicts undermine the overall value of the searchers’ bundles, despite the increase in their bid ratios. At this point, builders become more reliant on the value of their own bundles, and utilizing a lower rebate ratio allows them to retain a greater share of surplus. From Fig. 4, we can see that the probability of conflicts has a minimal impact on builders’ earnings. Conversely, the average reward for proposers exhibits a distinct pattern of initially increasing and then decreasing as the probability of conflicts rises, as shown in Fig. 4. An increasing probability of conflicts compels searchers to adopt higher bid ratios; however, excessively high probabilities of conflicts lead to a reduction in the overall MEV of the block. Under the current settings, a probability of conflicts $p_{C}$ of 0.5 maximizes MEV.
(a) Average bid ratio
(b) Average rebate ratio
(c) Average reward for searchers
(d) Average reward for builders
(e) Average reward for proposers
@@
III-C Empirical Game-Theoretic Analysis
In this subsection, we consider the agents’ selection between bundle sharing and block building strategies as two types of meta-strategies and conduct an empirical game-theoretic analysis. We consider a meta-game consisting of 10 agents, and we can obtain a heuristic payoff table $(N,U)$ [14], where each row $N_{i}$ contains a discrete distribution of $10$ players across two strategies. Here, $N_{i1}$ indicates the count of agents adopting block building, whereas $N_{i2}$ signifies the count of agents opting for bundle sharing. The right-hand side of the table $U_{i}$ represents the payoffs associated with the respective strategies, contingent upon the strategy profile $N_{i}$. By conducting repeated agent-based simulations for various strategy profiles $N_{i}$, we aggregate the corresponding strategy payoffs and compute the average as $U_{i}$.
TABLE I: Heuristic payoff table
$N_{i1}$ $N_{i2}$ $U_{i1}$ $U_{i2}$
1 9 $u_{11}$ $u_{12}$
2 8 $u_{21}$ $u_{22}$
1 2 $u_{31}$ $u_{32}$
9 1 $u_{91}$ $u_{92}$
Based on the heuristic payoff table, we employ the $\alpha$-Rank algorithm [15] to establish the Markov transition matrix for the two categories of strategies and calculate the stationary distribution of this transition matrix, which serves as a a dynamic solution for the meta-game to evaluate and rank the meta-strategies. Furthermore, we calculate the stationary distribution under varying probabilities of conflict and conduct a sweep over the ranking intensity alpha as noted in [15], ranging from $0.1$ to $100$. An adequately large ranking intensity guarantees that $\alpha$-Rank maintains the ranking of strategies most closely aligned with the Markov-Conley chains solution concept. Fig. 5 clearly demonstrates that an increase in the probability of conflicts leads agents to prefer the block building strategy, which entails building blocks to engage in competition within the block building auction, instead of resorting to the bundle sharing strategy where they share their bundles with builders to evade fierce competition. This result aligns well with intuitive reasoning. In particular, we can identify the critical probability of conflicts at which agents’ strategies transition, which is approximately $0.2$.

### v2

@@ III-B Modeling of Co-evolution
In our modeling, we employ genetic algorithms to capture the co-evolution process of agents. Each agent begins with a randomly generated population of 20 strategies. The block building strategy comprises a 5-digit binary sequence referred to as the “chromosome”, indicating the relevant rebate ratio $\alpha_{j}$. As shown in Fig. 2, for bundle sharing, $S_{i,r,k}$ signifies the $k$th strategy of searcher $i$ in the $r$th generation. Each 5-digit segment within a chromosome corresponds to a parameter value that regulates the bidding behaviors of bundle sharing strategy. The bid ratio is computed using a modified sigmoid function
@@
where $\gamma_{i,1}\in[1,5]$ measures the sensitivity of searchers’ bidding inclinations towards builders with elevated rebate ratio, whereas $\gamma_{i,2}\in[0,4]$ determines the overall scale of the bid ratio. The adoption of the modified sigmoid function ensures that the bid ratio $\beta_{ij}$ remains within the interval $[0,1]$ and is monotonically increasing in relation to $\alpha_{j}$. The largest decimal number that a 5-digit binary number can represent is $31$. Taking $S_{i,r,1}$ “0010101001” as an example, the first 5-digit binary segment “00101” can be translated to the decimal number number $51\cdot 2^{2}+1\cdot 2^{0}=$ 5. Thus, the parameter $\gamma_{i,1}$ can be derived from “00101” as $1+(5/31)\cdot(5-1)=21/31$. Similarly, the second 5-digit binary “01001” corresponds to the parameter $\gamma_{i,2}=0+(9/31)\cdot(4-0)=36/31$ suggesting that $S_{i,r,1}$ “0010101001” corresponds to the parameter pair $(\gamma_{i,1},\gamma_{i,2})=(21/31,36/31)$. The rebate ratio $\alpha_{j}$ for block building is directly obtained through a mapping that converts a 5-digit binary number into a decimal value within the interval $[0,1]$.
Figure 2: Searcher $i$’s strategy population in the $r$th generation
@@
Step 1: A private value is randomly generated for each agent, representing an opportunity available to them, alongside the random creation of an interaction matrix. In our subsequent experiments, the private values adhere to an exponential distribution with a default parameter of 10 (mean value of 0.1). [^4] Given a fixed mean (or rate parameter), the exponential distribution is the exclusive distribution that fulfills the maximum entropy requirement, thereby introducing the fewest extra assumptions. For simplicity, the interaction matrix is constructed according to a two-point distribution $\mathbb{P}(\varphi_{ij}=\varphi_{ji}=-1)=p_{C}$ and $\mathbb{P}(\varphi_{ij}=\varphi_{ji}=0)=1-p_{C}$. This implies that for any two agents, they completely conflict with a probability of $p_{C}$, or they operate independently with a probability of $1-p_{C}$. [^5] As pointed in [23], order flow providers possessing complementary flows are incentivized to integrate in order to collaboratively capture greater value, which consequently leads to a form of interaction between bundles that is predominantly characterized by conflict.
@@
$\mathbb{P}(S_{i,r,k}~{}\text{is selected})=\frac{e^{f(S_{i,r,k})/T}}{\sum_{k}e^{f(S_{i,r,k})/T}},$
where $T$ controls the trade-off between exploration and exploitation; through experimentation, we set $T=2$. At the beginning of each simulation, agents are randomly assigned a population of 20 strategies. Each agent maps its selected strategy (chromosome) to executable actions. Specifically, builder agents determine their rebate ratio $\alpha_{j}$, and searcher agents establish their bid ratio $\bm{\beta}_{i}$.
Step 3: Based on the individual strategies of the agents, the simulation of the block building process unfolds: searchers send bundles to builders, who utilize the bundles received from searchers along with their own bundles to construct the valuable blocks. We employed a greedy algorithm (Algorithm 1) to simulate the merging of bundles into a block. [^6] This study focuses on conditions of sufficient block space, wherein the total quantity of bundles does not exceed the capacity of the block. Nevertheless, our findings can be readily adjusted to address cases with capacity limitations on blocks, where a block can accommodate a maximum of $L$ bundles. Builders participate in the block building auction to calculate the final rewards received by both builders and searchers from the winning block. Update the strategy fitness based on the strategies selected by the agents in this round.
@@
Step 4: Agents optimize their strategy repository through genetic algorithms with a trigger probability of $0.01$. Initially, a subset of strategies is removed from the repository based on a specified elimination ratio (configured as $0.5$ in the simulation). Grounded in the “chromosomes”, genetic algorithms integrate selection, crossover, and mutation processes until the strategy repository is replenished back to its upper limit of $20$, as detailed below.
@@
  • Crossover: Two parent strategies are combined to create two offspring strategies by exchanging segments of the “chromosome” representations of the parent strategies. The fitness of offspring strategies is calculated as the mean of the fitness values of their parent strategies.
  • Mutation: For every bit in the binary chromosomes of each offspring produced, a flip ($0\rightarrow 1$ and $1\rightarrow 0$) is performed with a specified mutation rate of $0.01$.
IV Results
In this section, we present the results of agent-based simulations. We first analyze the co-evolution of agents using genetic algorithms, then assess how bundle interactions affect system dynamics and payoffs. Finally, we examine and predict participants’ role selection in the block building market using EGTA.
IV-A Co-evolution of Agents’ Strategies
(a) Evolution of CoV
@@ IV-A Co-evolution of Agents’ Strategies
In this subsection, we simulate a model with 10 builders, 10 searchers, and a conflict probability of $p_{C}=0.8$ over 10,000 blocks.
First, we assess the convergence of agent strategies by monitoring the coefficient of variation (CoV) for key strategy parameters over successive iterations. For each builder, 20 strategies for the rebate ratio $\alpha$ are maintained per iteration; each strategy, represented as a 5-bit binary encoding, corresponds to a discrete integer in $[0,31]$. The CoV for these values is computed as the standard deviation divided by the mean, with lower CoV values indicating greater consensus among strategies. Similarly, for searchers, we track the CoV of the behavioral parameters $\gamma_{1}$ and $\gamma_{2}$. Since all agents following the same meta-strategy are homogeneous, we report the average CoV for builders’ rebate ratios and for the searchers’ parameters across all agents in each iteration. Fig. 3(a) demonstrates a clear downward trend in the CoV of $\alpha$, $\gamma_{1}$, and $\gamma_{2}$, declining from $0.56$, $0.57$, $0.56$ to $0.12$, $0.24$, and $0.43$ over 10,000 iterations, respectively. These results indicate that agents’ strategies generally converge over time.
@@
To examine the evolutionary process of agents’ behaviors, we calculated the cumulative fitness of all strategies within the agents’ strategy repository for each iteration. Figs. 3(b), 3(c) and 3(d) depict the evolution of the cumulative fitness of $\gamma_{1}$, $\gamma_{2}$ and $\alpha$, respectively. The fitness reflects the performance of the strategy and the agents’ propensity to utilize it. On one hand, we observe that only a small subset of strategies achieve high fitness (lighter shades in the figure), clustering around certain values, while most strategies have near-zero fitness, consistent with the overall decline in CoV. However, these high-fitness strategies are not entirely concentrated, indicating some diversity in agent behaviors. Overall, searchers display differentiated bidding: they tend to offer higher bid ratios to builders with higher announced rebate ratios (Fig. 3(b)). They also generally choose smaller values of $\gamma_{2}$ to increase their overall bid ratios, reflecting intensified competition under $p_{C}=0.8$ (Fig. 3(c)). For builders, Fig. 3(d) shows that several extremely high rebate ratios correspond to high fitness, suggesting fierce competition and relatively low profitability for builders. We analyze agent profitability in detail in the following subsections.
@@
To provide a clearer illustration of the overall evolution of rebate ratios and bid ratios among the two types of agents within the system, we calculated the average values of bid ratios (for searchers) and rebate ratios (for builders) based on the actual actions taken in each iteration, as depicted in Figs. 3(e) and 3(f), respectively. The calculation of the moving average reveals a sustained upward trend in the average bid ratio, as shown in Fig. 3(e), alongside the convergence of the average rebate ratio toward approximately 0.7, as indicated in Fig. 3(f).
Overall, our agent-based simulation demonstrates that agents successfully adapt to the PBS environment by autonomously learning strategies through reinforcement learning. This co-evolutionary process leads to the convergence of strategies, while preserving some degree of diversity.
IV-B Impact of Bundle Interactions
@@
In this subsection, we focus on the impact of bundle interactions on the evolutionary outcomes of the system. During the simulation process, we simplify the interactions among bundles to include conflicts that arise between any two bundles with a specified probability. By performing 10 repetitions for each designated conflict probability within a system comprising 5 builders and 5 searchers, we statistically analyze the average bid ratio of searchers, the average rebate ratio of builders, and the profits of searchers and builders’ proposers after the system reaches a state of stability, yielding the results depicted in Fig. 4.
@@
According to Fig. 4, it is evident that the average bid ratio and average reward results of searchers exhibit less volatility compared to those of builders. For searchers, an increase in the probability of conflicts significantly intensifies competition among them, elevating the overall level of their bid ratios while concurrently reducing their average returns, as depicted in Figs. 4 and 4. The impact of the probability of conflicts on builders is relatively minor; however, it exhibits a more complex non-monotonic pattern. When the probability of conflicts is low, an increase in this probability encourages builders to raise their rebate ratios, a result of co-evolution: searchers correspondingly increase their bid ratios, thereby providing more valuable bundles to builders. In turn, builders enhance their rebate ratios to capitalize on the differentiated order flow from searchers. However, once the probability of conflicts surpasses a certain threshold (0.3 in Fig. 4), more frequent bundle conflicts undermine the overall value of the searchers’ bundles, despite the increase in their bid ratios. At this point, builders become more reliant on the value of their own bundles, and utilizing a lower rebate ratio allows them to retain a greater share of surplus. From Fig. 4, we can see that the probability of conflicts has a minimal impact on builders’ earnings. Conversely, the average reward for proposers exhibits a distinct pattern of initially increasing and then decreasing as the probability of conflicts rises, as shown in Fig. 4. An increasing probability of conflicts compels searchers to adopt higher bid ratios; however, excessively high probabilities of conflicts lead to a reduction in the overall MEV of the block. Under the current settings, a probability of conflicts $p_{C}$ of 0.3 maximizes MEV.
IV-C Empirical Game-Theoretic Analysis
In this subsection, we treat agents’ choices between bundle sharing and block building as two meta-strategies and analyze them using EGTA. We construct a meta-game with 10 agents, generating a heuristic payoff table $(N,U)$ [14], where $N_{i}$ specifies the distribution of agents across the two strategies ($N_{i1}$ and $N_{i2}$ represent the numbers of agents adopting block building and bundle sharing, respectively), and $U_{i}$ denotes the corresponding average payoffs derived from repeated agent-based simulations. Based on the heuristic payoff table, we employ the $\alpha$-Rank algorithm [15] to establish the Markov transition matrix for the two categories of strategies and calculate the stationary distribution of this transition matrix, which serves as a a dynamic solution for the meta-game to evaluate and rank the meta-strategies. Furthermore, we calculate the stationary distribution under varying probabilities of conflict and conduct a sweep over the ranking intensity alpha as noted in [15], ranging from $0.1$ to $100$. An adequately large ranking intensity guarantees that $\alpha$-Rank maintains the ranking of strategies most closely aligned with the Markov-Conley chains solution concept.